4. Which of the following statements is true, holding all else equal? I. If inflation increases and investors real rate of return stays the same, bond prices tend to decrease II. If investors real rate of return decreases and inflation stays the same, bond prices tend to increase III. If the yield to maturity on a bond increases, the bonds coupon rate will increase I onl II only 1 and 11 only II and III only C. e. I, II, and III 5. Assume a firm has sales of $2,750 on assets totaling $1,500, net income of $100, and dividends of $30. What is the sustainable growth rate if the equity has a value of $500? 20.0% 6.4% C. 12.7% -16.3%
Answer #1

Answer:

(4). Option B is the answer that is II only correct answer

If inflation increases ,the real rate return is decreased and bond prices will go up,so statement I is wrong.

If the yield to maturity increases or decrease , coupon rate will be same , but current yield are fluctuated. So option III also wrong.

Investor real rate of return decrease where inflation is stable then bond prices will go up. So option II is true.

(5). Sustainable growth rate= ROE×(1-dividend payout ratio)

ROE=net income ÷ equity

=$100÷$500=0.2

Dividend payout ratio =dividend÷net income

=$30÷$100=0.3

Sustainable growth rate =0.2×(1-0.3)

=0.2×0.7=0.14=14%